OpenStudy (anonymous):

3(3x - 1) + 2(3 - x) = 0 solve for x

OpenStudy (anonymous):

first u distribute: 9x-3+6-2x=0 then u combine like terms: 7x+3=0 move the 3 to the other side: 7x=-3 divide by 7 to get x by itself: x= -3/7 hope that helped :)

OpenStudy (anonymous):

Thank you, it did help!

OpenStudy (anonymous):

ur welcome :)